OpenStudy (anonymous):

The local minimum of the cubic function: y= -1/3x^3 + 6x^2 - 27x - 9 is found when x=?

OpenStudy (anonymous):

-x^2 + 12x - 27 = 0 -x^2+12x = 27 -x^2+x = 27/-12 x=-2.25

OpenStudy (anonymous):

correct or wrong?

hartnn (hartnn):

-x^2+12x = 27 x^2-12x+27=0 this is the quadratic equation you need to solve, you will get 2 roots of x.